• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name emily

    Log entry time 16:34:02 on July 29, 2006

    Entry number 176017

    This entry is a followup to: 176011

    keyword=BCM Calibration

    The BCM was calibrated to the 0L02 around 4pm today. We have data at all of 20uA,15uA,10uA,5uA and although it is not analyzed yet, we do have data that looks good to a first look. There were no beam trips this time :)